Shift summary, Day shift, Dec. 29, 1999

Vivian O'Dell (odell@fnal.gov)
Thu, 30 Dec 1999 06:52:56 -0600

Shift Summary for DAY shift, 29/Dec/99

On Shift: Sasha Glazov, Viv

Hours of beam delivered at all intensities: 7.0/8.5
Hours of beam delivered at nominal intensity: 7.0
Hours of E799 data (special runs not included): 6.7
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: End/Start run - 10 min, pipeline reset - 10 min,
cryo problems - 30 min, booster problems - 50 mins,
MI RF problems - 10 min

========================================================================
Runs ended this shift:
Run 15315 - BEAM_799, 19.0M.

Tapes brought to FCC this shift:
UPA280-282
UPB280-282
UPC280-282
UPS094

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Pretty darn smooth E799 datataking @ close to 1E13

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15315 07:00 New crackerjack shift
15315 last 09:09 beam is gone. cryo problems. 30 minute estimate.
We use this time to end run, swap tapes and
drt_fresh_start.
15316 09:35 Failed to start new run because tape drive on
ktev3 needs cleaning. We obediently clean it.
09:38 Beam returns 1 minute early. We're still cleaning
tape drive.
15317 09:42 start new run, finally.
15317 10:42 ktevana errors. Lose 1 spill.
15317 11:46 Booster problems. 15 minute estimate.
15317 12:35 Estimate proves crappy. Beam now back.
15317 14:00 Dave N. switches helium tubes to #'s 3 & 5.
15317 205 14:25 pipeline asks to print some file which
Sasha does. Man taking orders from machine.
The file is in the logbook. Pipeline did not
ask to look at it. Sasha resets pipeline system. Crate 11 didn't respond and got booted, I mean
rebooted.
15317 211 14:35 Running again.
15317 232 15:03 Main Injector RF problems. 10 minute estimate.
15317 238 15:14 Beam back. Meanwhile Jason L notices that
damp has stopped updating. We look at the
complicated instructions about what to do that
Margaret left in the control room. Viv votes to
wait 15 minutes so that the next shift can take
care of it. Unfortunately a member of the next shift
were also here and voted the motion down. Sasha
attacks the problem with vim and vigor.
15317 247 15:25 Sasha defibrillates the patient. As he strokes his
beard, he pronounces the heartbeat is now normal.
He is a true expert and we will replace Margaret's
pager number with Sasha's. Way to go, Sasha!
15317 251 15:32 Day shift passes off, with Sasha mumbling in
Russian something about writing the shift summary
next time.
========================================================================
Follow up Notes on Previous Shift
-Did a drt_fresh_start when the run ended.

========================================================================
Items to Follow up on Next Shift

-keep an eye (ear?) on the fan in rack 161. Replace if down time.

-page Sasha Golossanov if there is a 1+ hour shutdown (not owl shift).

- NOTE 3Y-6 KQ module is not being read out - so monitoring histogram
has a hole in it. This also has an effect on YTF.
This is a known readout problem.

=========================================================================